OpenStudy (anonymous):

(06.02)A bag contains 40 marbles, 4 of which are blue, 10 are red, 25 are green, and 1 is purple. Shawna takes a marble out of the bag, records the color, and returns it to the bag. How many green marbles should she expect after 400 trials? 25 40 100 250

OpenStudy (anonymous):

can anybody help?

OpenStudy (welshfella):

there are 25 green marbles out of a total of 40 so the property of picking a green marble is 25/40 = 5/8

OpenStudy (welshfella):

so the expected number in 400 tyrials is 400 * 5/8

OpenStudy (anonymous):

but thats not one of the answers

OpenStudy (welshfella):

work it out

OpenStudy (anonymous):

all it gives me is: 25 40 100 250

OpenStudy (welshfella):

multiply 400 by 5 then divide by 8

OpenStudy (anonymous):

the answers 200?

OpenStudy (welshfella):

no

OpenStudy (welshfella):

its just calculator work

OpenStudy (anonymous):

oh its 250 i was trying to do it in my head lol thank you so much :)

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

can you help me with another one

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

(06.02)A six-sided number cube labeled 1 through 6 is rolled 500 times. An odd number is rolled 325 times. Compare the experimental probability of rolling an odd number with the theoretical probability of rolling an odd number and select one of the statements below that best describes the situation. The experimental probability and theoretical probability are the same The experimental probability is larger than the theoretical probability The experimental probability is smaller than the theoretical probability There is not enough information to determine the relative frequency

OpenStudy (anonymous):

sorry it took me a while t o type out lol

OpenStudy (welshfella):

the experimental P of rolling odd = 325/500 = 13/20

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

what does that mean lol

OpenStudy (welshfella):

now what do you think the theoretical probability of getting an odd number is? Note there are an equal number of odd and even numbers on the cube

OpenStudy (anonymous):

oh A

OpenStudy (anonymous):

thats what i got before lol but i thought it was wrong xD

OpenStudy (welshfella):

the experimental probability is the result of experimenting - the 500 rolls of the cube.

OpenStudy (anonymous):

so A?

OpenStudy (welshfella):

no they are not the same the theoretical is what you might expect to get - due to the fact that there are an equal number of odd and even numbers

OpenStudy (anonymous):

im so confused

OpenStudy (welshfella):

on a coin you have 2 things heads or tails - what is Probability of getting a head?

OpenStudy (anonymous):

50%

OpenStudy (welshfella):

yes or as a fraction its 1/2 - well its same for odd numbers on a dice - 2 possiblities odd or even

OpenStudy (anonymous):

can you give me the answer lol i dont get any of this

OpenStudy (welshfella):

experimental probability = 13/20 theoretical probability = 1/2 so now you can pick the correct option

OpenStudy (welshfella):

- which is the largest?

OpenStudy (anonymous):

epirement

OpenStudy (welshfella):

correct Now you have it.

OpenStudy (anonymous):

thanks again

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

i wish you could help me with my whole test lol T^T

OpenStudy (anonymous):

im really stupid xD

OpenStudy (welshfella):

we are here to help you understand read over the posts and try to understand why we have done what we've done.

OpenStudy (welshfella):

you understand the coin part so you should understand the cube part you have 6 numbers on a cube 135 and 246 - that is 2 lots - one odd and one even so P(even) = 3/6 = 1/2 and P(odd) = 3/6 = 1/2 also

OpenStudy (anonymous):

but could you help me with the rest of the test, sorry to ask but i have 13 more mintues until my times up and then i wont be able to enter the tests again so im trying to rush

OpenStudy (anonymous):

and oh ok

OpenStudy (anonymous):

yeah i understand it better now

OpenStudy (welshfella):

no i haven't got the time for that . And thats not what openstudy is about

OpenStudy (anonymous):

ok but thanks for the previus help!